OpenStudy (anonymous):

w=2x+5y+6z . solve for y

OpenStudy (anonymous):

move all y terms to one side and all others to the oter side

OpenStudy (anonymous):

so w -2x - 6z = 5y right?

OpenStudy (anonymous):

Hi again... this is similar to your last, but you have to subtract first, then divide by 5 to solve for the y... give it a try and see

OpenStudy (anonymous):

@Nameless is right...

OpenStudy (anonymous):

ahh not just trying to help. ill step out.

OpenStudy (anonymous):

Yeaah @Nameless

OpenStudy (anonymous):

no worries, you just typed faster :)

OpenStudy (anonymous):

Ohh Goodness Thank You . I Got ; w=-6z-2x=5y ?

OpenStudy (shaik0124):

w -2x - 6z/5 = y

OpenStudy (anonymous):

careful... you have divide EVERYTHING on the left by 5... make sure you don't just divide the 6z term by 5... I think that's what @shaik0124 meant though

OpenStudy (shaik0124):

yup i divided everything by 5 but is is nt possible to type byy pc so it look like 6z/5

OpenStudy (anonymous):

Huh ? Wait Slowwww Down lol .All I Got right Now Is W=-6z-2x=5y . Now I Have subract or didvide or whatever by 5 ? what about the 2 ?

OpenStudy (anonymous):

* W-6z-2x=5y

OpenStudy (shaik0124):

here y will have 2 answers i.e,w/5=y this 1st answer and -6z-2x/5=y

OpenStudy (shaik0124):

@Isabel.A

OpenStudy (anonymous):

W-6z-2x/5=y ?

OpenStudy (anonymous):

w=2x+5y+6z . solve for y w = 2x + 5y + 6z w - 2x - 6z = (2x + 5y + 6z) - 2x - 6z w - 2z - 6z = 5y (w - 2z - 6z)/5 = 5y/5 = y

OpenStudy (shaik0124):

All I Got right Now Is W=-6z-2x=5y according to this it will have two answers else it will have 1 answer as i replied above i.e,W-6z-2x/5=y

OpenStudy (anonymous):

I don't think there is a reason for two answers... you just isolate y in terms of all the other variables

OpenStudy (anonymous):

Oh . lol . So W-6z-2x/5=y Is Right ?

OpenStudy (anonymous):

yes, as long as that /5 is actually dividing ALL the other stuff on the left

OpenStudy (shaik0124):

i said this answer earlier itself but again u said W=-6z-2x=5y . in that case i said y has 2 answers now decide which is the question n according to that choose the answer

OpenStudy (anonymous):

Ok v.v I Think I Get It Now . Lol . Thank You Guys Soooo Much !! It Soo Appericated .(: